The formula to find the sum of the first n whole number is
sum = (n^2 + n)/2
In order to encourage students to show up early to a school dance, the dance committee decides to charge less for those who come to the dance early. Their plan is to charge the first student to arrive 1 penny. The second student through the door is charged 2 pennies; the third student through the door is charged 3 pennies, and so on. How much money, in total, would be paid by the first 150 students?

Answer by richard1234(7193)

Pretty cheap school dance, even for 600 students.

1 + 2 + ... + 150 = 150(151)/2 = 11325. This is the number of cents, so the dollar amount is $113.25.